OpenStudy (anonymous):

square root(18)-sqrt(8) would that equal -sqrt(144) or sqrt(-144)

OpenStudy (anonymous):

\[-4*\sqrt{18}-\sqrt{8}\] this is what i'm trying to simplify.

OpenStudy (anonymous):

Neither? \[\sqrt{18} = \sqrt{9\cdot 2} = \sqrt{3^3 \cdot 2} = 3\sqrt{2}\] \[\sqrt{8} = \sqrt{4\cdot 2} = \sqrt{2^2 \cdot 2} = 2 \sqrt{2}\] \[\implies \sqrt{18} - \sqrt{8} = 3\sqrt{2} - 2\sqrt{2} = \sqrt{2}\]

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

Well then it's a bit different, but you can still use what I wrote to figure it out.

OpenStudy (anonymous):

Oh, except that the first part should read: \[\sqrt{3^2 \cdot 2}\] not \[\sqrt{3^3 \cdot 2}\]

OpenStudy (anonymous):

So if \[\sqrt{18} = 3\sqrt{2}\] Then \[-4\sqrt{18} = (-4\cdot 3)\sqrt{2}\]

OpenStudy (anonymous):

etc.

OpenStudy (anonymous):

this would end up to be \[-12\sqrt{2}-2\sqrt{2} \] subtract these the answers is \[-14\sqrt{2}\] that seem right.

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

cool thanks. appreciate your dedication to this site Polpak.